Two automobile dealer, A and B, had the same car which was listed at $18,000. Dealer A offered a $1200 rebate and would negotiate a further 8% discount after rebate. Dealer B offered an 8% discount and would negotiate a further $1200 rebate.
(a) Which dealer has the lowest price?
(b) What is the lowest price?(b

Answer by jorel1380(3719)

We should solve for both A & B's ending prices. Dealer A offered a $1200 rebate, plus an 8% discount after rebate,so:
A's price = (1-.08)(18000-1200) = $15456
Likewise, for dealer B we get:
(1-.08)(18000)-1200 = 16560-1200 = $15360
Thus the lowest rate is given by
(a) dealer B
at a rate of
(b) $15360